# OmniDRCA: Parallel Speech-Text Foundation Model via Dual-Resolution Speech Representations and Contrastive Alignment

### v1

@@ Appendix B Dataset Scaling
To investigate the impact of data scaling, we conduct experiments by progressively expanding the S2S training data on OmniDRCA-Small (w/o. Cont). Each S2S instance can be augmented into 7 distinct multimodal interaction patterns. As illustrated in Figure 3, expanding the dataset from 3.6K to 17.8K samples yields consistent improvements. The performance of T2M(S) exhibits nearly linear growth, while other patterns, despite showing slower growth rates, still demonstrate upward trends in the figure, suggesting potential for further performance enhancement through additional data.
Appendix C Prompt Templates
System prompts for multimodal interaction patterns.
@@ Appendix C Prompt Templates
Evaluation Templates.
To ensure fair comparisons given that different systems employ distinct data subsets for evaluation, datasets from two established evaluation frameworks were adopted: UltraEval-Audio [^7] https://github.com/OpenBMB/UltraEval-Audio for Speech-to-Speech ($S\rightarrow S$) evaluation and OpenAudioBench [^8] https://huggingface.co/datasets/baichuan-inc/OpenAudioBench for Speech-To-Text ($S\rightarrow T$) evaluation. For UltraEval-Audio, string-matching algorithms were applied to Llama Questions, Web Questions, and TriviaQA, while AlpacaEval utilized GPT-4o-mini with the Alpaca scoring template. In OpenAudioBench, GPT-4o-2024-08-06 was employed for scoring with different templates shown in Figures, all in strict compliance with the tools’ prescribed methodologies.
AlpacaEval Evaluation Template [Instruction] Please act as an impartial judge and evaluate the quality of the response provided by an AI assistant to the user question displayed below. Your evaluation should consider factors such as the helpfulness, relevance, accuracy, depth, creativity, and level of detail of the response. Begin your evaluation by providing a short explanation. Be as objective as possible. After providing your explanation, you must rate the response on a scale of 1 to 10 by strictly following this format: "[[rating]]", for example: "Rating: [[5]]". [Question] (Instruction Transcription) [The Start of Assistant’s Answer] (Response Transcription) [The End of Assistant’s Answer]
TriviaQA and Web Questions Evaluation Template Your will be given a question, the reference answers to that question, and an answer to be judged. Your tasks is to judge whether the answer to be judged is correct, given the question and reference answers. An answer considered correct expresses or contains the same meaning as at least **one of** the reference answers. The format and the tone of the response does not matter. You should respond in JSON format. First provide a one-sentence concise analysis for the judgement in field ‘analysis’, then your judgment in field ‘judgment’. For example, “‘json "analysis": "<a one-sentence concise analysis for the judgement>", "judgment": < your final judgment, "correct" or "incorrect"> ”’ # Question (instruction) # Reference Answer (targets) # Answer To Be Judged (answer_to_be_judged)
Llama Questions Evaluation Template ## Background You are a professional QA evaluation expert. You need to assess whether the model’s answer is correct based on the standard answer. ## Scoring Criteria Correct: The answer matches or is equivalent to the standard answer Incorrect: The answer is wrong or irrelevant to the question ## Evaluation Guidelines 1. The expression of answers can be flexible, not requiring exact matches. For example: - Numbers can be expressed in either Arabic numerals or words - Proper nouns can be in either English or Chinese - Differences in punctuation can be ignored 2. Focus on whether the core meaning of the answer is correct ## Output Format Provide the reasoning for your score, then generate the result in "[]" format and make sure it contains "the score is [Correct]" or "the score is [Incorrect]", for example: “‘ The answer is correct and equivalent to the standard answer, the score is [Correct] ”’ or “‘ The answer is incorrect and does not match the standard answer, the score is [Incorrect] ”’ ## Question: (instruction) ## Standard Answer: (targets) ## Model’s Answer: (answer_to_be_judged)

### v2

@@ 4.2 Results and Discussions
To investigate the impact of data scaling, we conduct experiments by progressively expanding the S2S training data on DrVoice-Small (w/o. Continous Speech Encoder). Each S2S instance can be augmented into 7 distinct multimodal interaction patterns. As illustrated in Figure 3, expanding the dataset from 3.6K to 17.8K samples yields consistent improvements. The performance of T2M(S) exhibits nearly linear growth, while other patterns, despite showing slower growth rates, still demonstrate upward trends in the figure, suggesting potential for further performance enhancement through additional data.
5 Conclusions and Future Work
We introduce DrVoice, a novel parallel speech-text voice conversation model that leverages joint autoregressive modeling with dual-resolution speech representations. Experimental results demonstrate that DrVoice sets a new state-of-the-art (SOTA) benchmark in both speech-to-speech and speech-to-text generation tasks against current baselines. Notably, its dual-resolution mechanism significantly improves inference speed and computational efficiency without sacrificing performance.
Our future work will address three key areas: (1) Enhancing Speech Generation Quality: To overcome the limitations of the current suboptimal tokenization and further improve speech fidelity, we plan to explore advanced speech tokenizers, such as the one utilized in CosyVoice. This will enable the model to capture finer acoustic details and generate more natural-sounding audio. (2) Enabling Full-Duplex Interaction: A crucial direction is to enable full-duplex interaction for more natural conversations. Inspired by recent advancements like Parrot (Wang et al., 2025), we will investigate the use of a time-division multiplexing (TDM) input stream. This would permit DrVoice to accept user speech inputs during its own speech generation phase, thereby efficiently utilizing idle time slots and allowing for responsive, interruptible dialogue. (3) Expanding to General Audio and Multimodality: Finally, we aim to extend our research beyond speech-centric tasks to broader audio comprehension and synthesis, including the recognition and generation of music and environmental sounds. The subsequent integration of the visual modality will be a critical step toward developing a more comprehensive, multimodal conversational AI.
Appendix A Prompt Templates
@@
Table 6: System Prompts for Multimodal Interaction.